-4(k-7)-(-5k-7)=-8
Answer by ptaylor(2198)

-4(k-7)-(-5k-7)=8 get rid of parens
-4k+28+5k+7=8 collect like terms
-4k+5k+28+7=8
k+35=8 subtract 35 from each side
k=-35+8
k=-27
CK
-4(-27-7)-((-5*-27)-7)=8
-4*(-34)-(135-7)=8
136-128=8
8=8
Hope this helps---ptaylor